Gucchi:

Solve negative 5 over 2, the whole multiplied by x minus 5 equals negative 55. โˆ’20 โˆ’24 20 24

Gucchi:

I picked 2nd one just wanna make sure its right

Gucchi:

its either 2nd or third idk

supie:

Why the 3rd one?

Gucchi:

i'll show you hold on

Gucchi:

-(5/2)x -5 = -55 -(5/2)x -5 + 5 = -55 + 5 -(5/2)x = -50 -(5/2)x ร— 2 = -50 ร— 2 -5x = -100 x= -100 รท -5 x= 20

Gucchi:

but then i looked at the question and saw that it said minus 5 and not negative 5

supie:

This is my work... First we write out the equation \[โˆ’5/2xโˆ’5=โˆ’55\] +5 to both sides \[โˆ’5/2xโˆ’5+5=โˆ’55+5\]\[โˆ’5/2x=โˆ’50\] x both sides by 2/(-5) \[(2โˆ’/5)*(โˆ’5/2x)=(2/-5)*(โˆ’50)\]

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gucchi but then i looked at the question and saw that it said minus 5 and not negative 5 \(\color{#0cbb34}{\text{End of Quote}}\) So yeah I think you're correct about option 3.

Gucchi:

does that equation lead you to option 3?

supie:

Yes.

supie:

You're correct about option 3 :)

Gucchi:

I found this guy online doing it a different way and he came up with a different answer. Can i copy and paste it here?

Gucchi:

I was just confused by that.

supie:

Oh sure.

Gucchi:

actually i'll send pic

Gucchi:

Gucchi:

i need to do something quickly brb

Gucchi:

i'll keep question open. ill be back in like 5 mins

supie:

Ok. I think that might have been a different problem or something because I think thats wrong.

Gucchi:

Yeah, its probably wrong, because he wasn't adding both sides like how you were doing it

supie:

Yeah, so the correct answer has to be the 3rd option Right?

Gucchi:

i'll stay with your answer, thanks for your help!

Gucchi:

yh it has to be

Gucchi:

i have another question here thats a little confusing

Gucchi:

Which statement is true for the equation 4x โˆ’ 4x โˆ’ 5 = โˆ’5? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

Gucchi:

I think its the last one

supie:

Ok so, \[4xโˆ’4xโˆ’5=โˆ’5\] Simplify.... \[4xโˆ’4xโˆ’5=โˆ’5\]\[4x+โˆ’4x+โˆ’5=โˆ’5\] Combine the like terms \[(4x+โˆ’4x)+(โˆ’5)=โˆ’5\]\[โˆ’5=โˆ’5\] \[โˆ’5+5=โˆ’5+5\] So what do you think the answer is?

Gucchi:

the last one because 4xs cancel themselves out

Gucchi:

your left with -5=-5

supie:

Yeah, I think so too. Good luck! :)

Gucchi:

Thank you!

supie:

Np!

Gucchi:

Alec calculated the slope of line AC between points A and B. Avis calculated the slope of line AC between points B and C. Both Alec and Avis drew triangles, as shown below, to calculate the slopes: A graph shows numbers from 0 to 18 on the x axis at increments of 2 and the numbers 0 to 6 on the y axis at increments of 1. A line AC is drawn that joins the ordered pairs 0, 0 and 18, 6.The ordered pair 6, 2 is labeled as B. A right triangle is drawn with AB as hypotenuse and height 2 units. Another right triangle is drawn with BC as the hypotenuse and height 4 units. Which statement is correct about the slopes calculated by Alec and Avis if both calculated the slopes accurately? Both calculated the slope as 3. Both calculated the slope as 1 over 3. Alec calculated the slope as 1 over 3 and Avis calculated it as 3. Alec calculated the slope as 3 and Avis calculated it as 1 over 3. I'll include pic if i can

Gucchi:

Gucchi:

I know how to calculate slopes, but i'm confused here

Gucchi:

do i get the slope of the first triangle then simplify it?

Gucchi:

and do the same with the other one?

Gucchi:

i would get 3 for both

supie:

To find slope you have to you divide the difference of the y-coordinates from 2 points on a line by the difference of the x-coordinates of those same 2 points

Gucchi:

which two points do i have to divide here?

supie:

@jhonyy9

Gucchi:

i'm thinking its the second one

Gucchi:

Ricky needs $45 to buy a jacket. He has saved $15 and plans to work as a babysitter to earn $5 per hour. Which inequality shows the minimum number of hours, n, that Ricky should work as a babysitter to earn enough to buy the jacket? 5n โ‰ฅ 45 + 15, so n โ‰ฅ 12 5n โ‰ค 45 + 15, so n โ‰ค 12 15 + 5n โ‰ฅ 45, so n โ‰ฅ 6 15 + 5n โ‰ค 45, so n โ‰ค 6

Gucchi:

i picked third, wanted to make sure its right

Gucchi:

i got the slope one right it was the second one

supie:

Ok good job! This one is easy, what do you think? btw not D.

Gucchi:

I was thinking c?

supie:

Yeah same.

Gucchi:

If a sprinkler waters 1 over 18 of a lawn in 1 over 3 of an hour, how much time will it take to water the entire lawn? 6 hours 9 hours 18 hours 21 hours

Gucchi:

i picked a

supie:

I think ur right Sry i was afk

Gucchi:

It's alright

supie:

ok thanks. You got any more?

Gucchi:

Which unit rate is the lowest price per ounce? Choice A: 5 ounces of raisins for $1.49 Choice B: 12 ounces of raisins for $3.69 Choice A Choice B The unit rates are equal. The unit rates cannot be determined. This has to be a right?

Gucchi:

I have a few more

Gucchi:

@supie is this A?

Gucchi:

I divided 1.49 by 5

supie:

I think so...

Gucchi:

and 3.69 by 12

supie:

Yeah so, it has to be A. I agree.

Gucchi:

A theater company charges $10.50 per ticket for people of all age groups. If c represents the total number of children who bought tickets and a represents the total number of adults who bought tickets, which expression best represents the total money earned by the company from the sale of tickets? 10.5c + a c + 10.5a 10.5(c + a) 21(c + a)

Gucchi:

this is another one, i picked c

supie:

Yeah C sounds right...

supie:

its 10.50/person so yeah \[10.50 * (15 + 4)\]

Gucchi:

true

Gucchi:

ill pick that

Gucchi:

The price of tomatoes at a farm is $0.54 per pound. Which equation can be used to determine c, the total price of n pounds of tomatoes? c = 0.54n c = 0.54 over n c = 0.54 + n c = 0.54 โˆ’ n

Gucchi:

i picked the first one

Gucchi:

it makes the most sense right?

supie:

Yeah, the first one has to be correct.

Gucchi:

Which graph shows a proportional relationship between the number of hours of renting a costume and the total amount spent to rent the costume? There are 4 pics i will include.

Gucchi:

1 attachment
Gucchi:

1 attachment
Gucchi:

im thinking the last one

supie:

Maybe that's right but I'm not 100% sure

Gucchi:

@umm what do you think it is?